OpenStudy (anonymous):

Which of the following is a consumer durable? food kitchen appliances dishes DVDs

OpenStudy (anonymous):

consumer goods that are not destroyed by use

OpenStudy (anonymous):

Based on that definition it would stand to reason that food is consumed, kitchen appliances do not get destroyed after use, dishes do not get destroyed after use (except during a greek wedding), dvds do not get destroyed after use.

OpenStudy (mathstudent55):

Which question does not belong in the Mathematics section? a. 1 + 1 = ? b. 2x + 1 = 5, x = ? 3. a^2 = 16, a = ? 4. Which consumer goods are durable?

OpenStudy (anonymous):

Which two labels do not belong? 1.) a 2.) b 3.) 3 4.) 4

OpenStudy (mathstudent55):

LOL, you got me!!!

OpenStudy (anonymous):

That was pretty funny though.